The perimeter of a rectangle is 16 feet. The area of the rectangle is 15 square feet. the Length is greater then the width . What is the Length and the width?

Answer by mananth(16949)

let length be x
width =y
x*y=15
2x+2y=16
x+y=8
x=8-y
xy=15
y(8-y)=15
8y-y^2-15=0
y^2-8y+15=0
y^2-5y-3y+15=0
y(y-5)-3(y-5)=0
(y-3)(y-5)=0
y=3 , 5

the sides are 3 & 5 feet